OpenStudy (pottersheep):

Please help me with these gas problems (Grade 11) 1a) The kelvin temp of a 3L sample of oxygen gas is tripled while the pressure is help constant. What is the new volume of oxygen? b) Air trapped in a syringe at 100KPA of pressure has its volume decreased to 1/8 of its initial volume. If the temp is constant, the pressure in the syringe is?..

OpenStudy (anonymous):

ok so first start by writing ideal gas equation: pV=nRT now you have same amount of oxygen before and after so you can write next: n1=n2 and as n=pV/RT and R is constant so it would cross itself out you are left with: p1V1/T1 = p2V2/T2 so you have that pressure is constant so you can cross it out too and you are left with: V1/T1=V2/T2 hope you get it...

OpenStudy (pottersheep):

oo got the first one thanks! now ill work on the second one

OpenStudy (anonymous):

oh great, somebody who is willing to use head! glad to help

OpenStudy (pottersheep):

Yayy i got the second one too now :)

OpenStudy (anonymous):

congrats!

OpenStudy (pottersheep):

thanks! :)

OpenStudy (anonymous):

not a problem.